What is the molarity of a sodium carbonate solution containing 32.52 g Na2CO3 dissolved in enough water to make 822 ml of solution?


1
Expert's answer
2021-01-25T04:30:55-0500

According to the equation:

M = n / V = m / (Mr × V)

where M - molarity, n - number of moles, V - volume (V = 822 ml = 0.822 L), m - mass (m = 32.52 g), Mr - molecular weight (Mr(Na2CO3) = 106 g/mol).

From here:

M (Na2CO3) = 32.52 g / (106 g/mol × 0.822 L) = 0.373 mol/L


Answer: 0.373 mol/L